### v1

# No Size Fits All: The Perils and Pitfalls of Leveraging LLMs Vary with Company Size

Abstract
Large language models (LLMs) are playing a pivotal role in deploying strategic use cases across a range of organizations, from large pan-continental companies to emerging startups. The issues and challenges involved in the successful utilization of LLMs can vary significantly depending on the size of the organization. It is important to study and discuss these pertinent issues of LLM adaptation with a focus on the scale of the industrial concerns and brainstorm possible solutions and prospective directions. Such a study has not been prominently featured in the current research literature. In this study, we adopt a threefold strategy: first, we conduct a case study with industry practitioners to formulate the key research questions; second, we examine existing industrial publications to address these questions; and finally, we provide a practical guide for industries to utilize LLMs more efficiently.
No Size Fits All: The Perils and Pitfalls of Leveraging LLMs Vary with Company Size
Ashok Urlana1   Charaka Vinayak Kumar1   Bala Mallikarjunarao Garlapati1   Ajeet Kumar Singh1   Rahul Mishra2 TCS Research, Hyderabad, India1       IIIT Hyderabad2 ashok.urlana@tcs.com, charaka.v@tcs.com, balamallikarjuna.g@tcs.com ajeetk.singh1@tcs.com, rahul.mishra@iiit.ac.in
1 Introduction
Large language models (LLMs) have recently garnered significant attention due to their exceptional performance in various predictive and generative tasks Hadi et al. (2023); Kar et al. (2023). Extensive research has been conducted to harness LLMs across diverse domains and tasks Raiaan et al. (2024), including medicine Thirunavukarasu et al. (2023), finance Li et al. (2023b), and reasoning tasks Huang and Chang (2023); Qiao et al. (2023). Despite their unprecedented adaptation to numerous industrial applications, there is a notable lack of studies examining the potential challenges and risks associated with LLMs, which can vary depending on the size of the organization. Such studies would not only be valuable for industries seeking informed adaptation but also help shape research focus to address the key challenges and obstacles faced in real-world scenarios.
The challenges and bottlenecks faced by organizations of different sizes are not uniform. Factors such as funding availability, workforce size, skill and training deficits, ethical and regional considerations, and access to adequate hardware can all influence how these challenges manifest. Previous research has largely addressed general challenges Raiaan et al. (2024) with LLMs, such as multilingual support, domain adaptation, and compute requirements. However, there is a lack of studies specifically focusing on the industrial perspective and the unique challenges of implementing LLMs in this context.
To this end, we conduct a study with a threefold strategy, firstly, we conduct a rigorous case study of real-world practitioners from the IT industry, who are trying to work on AI adaptation and formulate three guiding research questions. RQ1. How have industries adopted LLMs so far, and what challenges do they face? RQ2. What are the barriers hindering the full utilization of LLMs in industrial applications, and how can these barriers be overcome? RQ3. How can various industries advance to maximize the utility of LLMs in practical applications? Subsequently, with an aim to address guiding research questions, we perform a thorough scoping survey of existing research publications from industrial entities of all sizes. Finally, we discuss our takeaways and insights and present a practical pilot scenario-based guide for industries to adapt to LLMs in a more informed manner.
The key contributions of this work can be summarised as: this study identifies various categories of challenges associated with LLMs for industrial adoption and proposes potential solutions. These challenges broadly relate to data confidentiality, reliability of LLM responses, infrastructure bottlenecks across industries, domain-specific adoption, synthetic data generation, and ethical concerns. Additionally, we offer a practical guide tailored for small, medium, and large industries to maximize the utilization of LLMs.
2 Related Work
In the literature, numerous studies focus on practical and ethical challenges associated with LLMs across diverse application domains includes education Yan et al. (2024), finance Li et al. (2023d), healthcare Zhou et al. (2023) and security Shao et al. (2024). Additionally, several studies address the task-specific challenges for LLMs’ adoption in areas such as spoken dialog systems Inoue (2023), mathematical reasoning Ahn et al. (2024), mining software repositories Abedu et al. (2024). Moreover, studies explore the challenges based on LLMs capabilities with explanations generation Kunz and Kuhlmann (2024), data augmentation Ding et al. (2024), support for multilingual context Shen et al. (2024) and compliance with ethical challenges Jiao et al. (2024).
Close to our work, Gallagher et al. (2024) addresses a few concerns on the adoption of LLMs for specific high-stake applications, particularly intelligence reporting workflows. In contrast to existing studies, our work specifically concentrates on the utilization of LLMs for industrial applications. Moreover, this study provides a comprehensive overview of several roadblocks to LLMs adoption for industrial use cases and corresponding potential solutions. Additionally, our study offers a suggestive guide to maximize the utilization of LLMs for various industries.
3 Methodology
This section aims to understand, how have industries adopted LLMs so far, and what challenges they face? (RQ1).
3.1 Industrial Case Study on LLMs
We conduct an industrial case study to understand, how the LLMs are shaping industry practices, identifying both challenges and benefits. Through a meticulous process of expert consultation and iterative refinement, the questionnaire was designed to capture insightful data and serve as a tool for understanding the evolving role of LLMs in the industry. This case study covers a multitude of aspects related to LLM usage for specific application domains, corresponding risks, trust attributes, and challenges. In crafting a succinct questionnaire, our objective was to gauge the adoption and impact of LLMs in various industries. These questions can be found in Appendix B Table 4. We receive 26 responses in total from real-world practitioners of the IT industry.
Refer to caption (a) Data modalities considered. Refer to caption (b) Foreseen risks associated with LLMs. Refer to caption (c) Opted trust attributes.
Figure 1: Industrial case study statistical overview of various aspects
3.2 Quantitative Analysis
Based on the responses obtained from the industrial case study, we make the following observations.
Participants of the case study. We shared the questionnaire with the IT professionals, who are either working on LLMs or developed some solutions. The participants are industry professionals and practitioners ranging in their expertise from beginner to expert level.
Widely adapted applications by leveraging LLMs. Even though LLMs are being utilized for various applications, we observe that the majority of the industrial applications are related to financial, retail, security, and healthcare domains.
Modality of the datasets. More than 60% of the industry practitioners prefer to use either textual or tabular data as shown in Figure 1.a.
Widely used LLMs. Our case study suggests that more than 50% of the applications utilize the GPT-3.5 and GPT-4 models. Recently, researchers have been assessing the capabilities of LLaMA-2 Touvron et al. (2023) and Mistral Jiang et al. (2023a).
Prompting strategy. We observe that zero-shot, few-shot, and in-context learning prompting strategies are widely adapted compared to Fine-tuning.
Risks associated with LLMs. Based on our case study, LLMs pose risks associated with security and safety, quality of service, and license-related challenges as depicted in Figure 1.b.
Trust attributes to be considered. We observe that robustness, security, and hallucination are the major attributes that need to be considered to utilize the LLMs as shown in Figure 1.c.
Moreover, to gain a better understanding of the barriers to leverage the LLMs for industrial use cases, we survey 68 research papers specifically from the industry. With this study, we compile several prominent challenges and collate potential solutions to address them. The selection criteria for the papers can be referred to in the Appendix A.
4 Challenges and Potential Solutions
In this section, we explore several barriers to leveraging LLMs for industrial applications and discuss potential solutions (RQ2).
4.1 Data Confidentiality
4.1.1 Pre-training data issues
Potential privacy risks. To deploy large language models (LLMs) on cloud platforms, robust data privacy protocols are required to handle extensive sensitive datasets while pre-training. Key challenges include mitigating data breaches and preventing unauthorized extraction of sensitive information. Despite the adoption of LLMs in applications like disaster response management Goecks and Waytowich (2023), public health intervention Jo et al. (2023), and assisting AAC users Valencia et al. (2023), there is a noticeable lack of focus on privacy and security aspects. Moreover, it is imperative to address the potential risks associated with deploying LLMs in high-stakes scenarios.
Regulations. GDPR in Europe and CCPA in California introduce stringent guidelines for deploying LLMs by enforcing strict data handling and intellectual property rules for transparency and fairness. As highlighted by Mesko and Topol (2023), adhering to these laws in sensitive domains like healthcare is crucial to avoid harm and protect privacy.
Potential solution. Developing a comprehensive framework that aids in LLM compliance is essential for responsible use and interaction with users.
4.1.2 Usage of APIs
To access the closed-source LLMs, passing the commercial data through third-party APIs raises potential privacy concerns Laskar et al. (2023).
Potential solutions. 1. Implementing robust security and privacy techniques, such as federated learning, is essential to safeguard user data while maintaining the functionality of LLMs, 2. A strategic way of crafting prompts is essential to avoid PII leakage Kim et al. (2023).
4.2 Reliability of LLMs’ Responses
Control the level of AI proactivity. LLMs should reduce social awkwardness, improve expressiveness, and adapt to different scenarios Liu et al. (2023c); Urlana et al. (2023). The open-ended generation of LLMs makes it challenging to customize dialog systems for public health intervention applications Jo et al. (2023).
Outdated knowledge. The open-endedness of LLMs often leads to hallucinations due to a lack of an updated knowledge base Faizullah et al. (2024). Additionally, the training data might contain errors and become outdated over time.
Potential solutions. Techniques such as Retrieval-Augmented Generation (RAG) are effective in reducing hallucinations. However, RAG systems struggle with complex questions requiring additional information and generate out-of-context content. Moreover, techniques such as diverse beam search Vijayakumar et al. (2016), confident decoding Tian et al. (2019) are promising in mitigating hallucinations. Additionally, model editing techniques Hoelscher-Obermaier et al. (2023) can address the unintended associations, enhancing the practical usage of LLMs.
4.3 Infrastructure Accessibility
Carbon emissions. Infrastructure is crucial for deploying LLMs, influencing factors like processing speed, latency, cost, and training needs. High-performance hardware is necessary to boost speed and reduce latency, enhancing user experience but demanding careful budgeting due to high costs. Achieving an optimal balance between cost and performance is crucial for the efficient training and scalability of LLM applications.
Potential Solution. Implementing robust small language models lead to reduced carbon emissions.
Compute requirements. Despite the state-of-the-art performance of the large language models, utilizing them for small-scale industries is not feasible due to high compute requirements.
API costs. While LLMs like GPT-3.5 and GPT-4 Achiam et al. (2023) demonstrate superior performance over open-source models, their high cost of API access is prohibitively expensive to perform comprehensive studies Laskar et al. (2023).
Potential Solution. Balancing the trade-off between performance and cost is necessary for the practical usage of LLMs Laskar et al. (2023).
High inference latency. APIs can be slow when demand is high. For instance, tasks like business meeting summarization can take GPT-4 around 40 seconds to generate a single response Laskar et al. (2023). Additionally, longer prompts increase computational demand Jiang et al. (2023b).
Potential Solution. Open-source models like LLaMA-2 Touvron et al. (2023) are more favorable for industrial deployment. Further studies on efficient model optimization techniques such as quantization, pruning, and distillation are required Laskar et al. (2023). Moreover, closed-source models that can utilize prompt compression techniques such as LLMLingua Jiang et al. (2023b) are promising.
4.4 Domain Adaption
Lack of domain-specific datasets. The ability of LLMs in the finance and medical domains is lacking due to insufficient Liu et al. (2023a) domain-specific training data in the foundation models Li et al. (2023c). Consequently, the current versions of GPT-4 and ChatGPT do not meet the industrial requirements to build financial analyst agents Li et al. (2023c). While LLMs can generate relevant reasoning, they fall short of the desired standard, indicating significant room for improvement.
Lack of expert domain data. Majority of the LLMs struggle to perform the expert data domain tasks Alfassy et al. (2022), such as image-text and text-image retrieval in technical documentation.
Diversity. Currently, LLMs fail to mitigate social bias due to the absence of diverse demographic groups Lee et al. (2023) data in the training data. Foundation models should give equal importance to aspects such as ethnicity, nationality, gender, and religion. At present, most LLMs reflect the collective intelligence of Western society.
In-context learning (ICL). The scope of in-context learning is limited by its pretraining data Han et al. (2023). It is unlikely that any model will perform well when using ICL with data significantly different from its pretraining data.
Potential Solution. 1. Pre-training data should consist of various domain mixtures; however, finding the right mixture is an open challenge. 2. LLMs should be carefully tested to ensure they treat marginalized individuals and communities equally Kotek et al. (2023). 3. Continuous pre-training can help overcome the drawbacks of the in-context learning strategy.
4.5 Data Creation Using LLMs
Few works attempt to generate synthetic datasets by utilizing LLMs. However, four major concerns exist with using LLMs for synthetic data creation/annotation; 1). Lack of diversity. Synthetic datasets may lack diversity due to the limited knowledge base Ramakrishna et al. (2023) of LLMs, 2). Quality and resource requirements. The quality of the annotated data might increase with the size of the LLM used for the annotation Sun et al. (2023). However, leveraging large LLMs requires higher computational resources, 3). In-context learning (ICL) challenges. ICL is a widely adopted approach for textual task data annotation tasks Li et al. (2023c). However, the challenge lies in responsibly incorporating the model’s output to deliver value to users without misleading them or inadvertently amplifying malicious behavior Deng et al. (2023), 4). Performance of complex tasks. The performance of models degrades when using synthetic data for complex tasks.
Potential solution. Currently, tools like FABRICATOR Golde et al. (2023), support tasks like classification, sentence similarity and QA for data labeling and other tasks should be explored.
4.6 Sub-standard Performance of LLMs
Code generation. LLMs’ coding ability is limited to generate general-purpose coding tasks. However, the generation of high-quality code for complex network management tasks remains challenging Mani et al. (2023). Moreover, LLMs have limited capabilities in repository-level coding tasks except in C and Python languages Bairi et al. (2023) and fail to complete code with potential bugs Dinh et al. (2023). Most of the code-LLMs struggle with code completion tasks, with undefined names and unused variables Ding et al. (2023) being the most prominent static error cases.
Conversational applications. LLMs face challenges in providing emotional support and maintaining long-term memory, impacting their effectiveness in conversational applications Jo et al. (2023). Future research on a longitudinal deployment of LLM-driven chatbots for public health interventions would help understand how users’ engagement changes over time.
Multilingual. Maximizing LLMs’ potential in healthcare, retail, and finance requires high-quality, labeled datasets. Additionally, with most LLMs being limited to English, there is significant room for creating robust multilingual models.
Multi-modal. With the rising demand for multi-modal applications Feng et al. (2023); Lu et al. (2023b), only a few studies focused on utilizing LLMs for such industrial applications. More efforts are needed to integrate LLMs with voice assistants and Robotics Yamazaki et al. (2023).
4.7 Evaluation of LLMs
In sectors like legal, finance, and healthcare, blending LLMs with human feedback is crucial to lowering false positives, underscoring the importance of human oversight in safety-critical applications Liu et al. (2023b). Moreover, our analysis (see Appendix C) reveals that less than 15% of studies conduct human evaluations to assess LLM outputs, indicating a need for more rigorous validation methods. Evaluating long-form question answering is challenging for LLMs Zhao et al. (2023), as additional contextual information may not always be available in practical QA scenarios. Current metrics like BLEU Papineni et al. (2002) and ROUGE Lin (2004) primarily evaluate the explanation task, but these measures are insufficient for assessing the reasonableness of LLM responses.
Small Scale Industries Medium Scale Industries Large Scale Industries
Mode of LLMs
usage
API Integration, Pre-trained Models,
Low-code or No code platforms,
Zero-shot, Few-shot
Domain-specific Fine-tuning,
Chain of thought
In-house deployment, Continuous pre-training,
Collaborative tools and frameworks,
Pre-training from scratch, Re-pretraining
Challenges
Cost, Technical expertise, Data
privacy, Performance
Scalability,
Domain Adoption
Ethical concerns, Regulations,
Data governance
Data modalities Uni-modal Multi-modal (Max two) Multi-modal (2 or more)
Training time Few hours to days Few days to weeks Few weeks to months
Dataset size 100 to 10k samples 10k to 100k samples More than 100k samples
Compute resources Cloud Cloud and Moderate GPUs In-house high-end GPUs and TPUs
Optimization Quantization
PEFT techniques,
Distillation, Pruning
Prompt compression techniques
Languages Monolingual Monolingual Multi-lingual/Cross-lingual
Ethical complexity Low Moderate to high High to very high
[…]
Table 1: A suggestive guide to various industries to maximize the utilization of LLMs for NLG applications.
4.8 Ethical Concerns
The most common ethical challenges with LLMs are violation of the model license, model theft, copyright infringement, producing harmful content, and trustworthiness Foley et al. (2023) and following are the potential solutions.
Protecting LLMs. Watermarking techniques Peng et al. (2023) are essential for copy-right protection of industrial LLMs, aim to minimize the adverse impact on the original LLM.
Building trust. LLMs face gender bias issues while generating professional documents like recommendation letters Li et al. (2023c). Further, most of the studies are limited to understanding gender bias in male and female pronouns, indicating a need for more research on marginalized groups, such as the LGBTQIA+ community. Moreover, AI-assisted writing should be employed judiciously to prevent reinforcing gender stereotypes Wan et al. (2023). Increasing trustworthiness involves prioritizing the addressing, opposing, and limiting of further proliferation of bias Kotek et al. (2023). Recent studies also suggest to conduct experiments with LLMs’ at least three times to validate the results Chen et al. (2023).
Enhancing creativity. The involvement of AI models should not aim to replace the author’s thinking but rather enhance it by producing new ideas and discover connections or findings Shen et al. (2023).
Fairness in data visualization. Developing interactive data visualization techniques/frameworks to identify the hidden biases and ensure fairness Kwon and Mihindukulasooriya (2023).
Linking models. Techniques such as LLM Attribution Foley et al. (2023) are promising for linking fine-tuned models with their pre-trained versions.
Protecting integrity. To maintain the integrity of LLMs, implementing guardrails such as NeMo Rebedea et al. (2023), LangKit [^1] https://docs.whylabs.ai/docs/langkit-api/, and TrustLLM Sun et al. (2024) helps to prevent the generation of biased, offensive, or inaccurate responses.
Figure 2: Current state of the industrial applications utilizing the LLMs; POC stands for proof of concept.
Addressing these challenges requires a combination of technical expertise, ethical considerations, and ongoing research efforts. As shown in Figure 2, we categorize each paper (total of 68) based on its application life cycle and observed that, due to the above-mentioned pitfalls, more than 70% of LLM-based studies are still in the conceptual phase.
5 Maximizing LLM Utilization Across Industries
This section offers a suggestive guide to various industries to maximize the utilization of LLMs for Natural Language Generation (NLG) applications (RQ3). As shown in Table 1, our suggestions are tailored to various industries, considering their distinct goals, resources, and workforce capabilities. Furthermore, the recommendations intended for small and medium-sized industries equally apply to large-scale industries.
1) Small-scale industries such as startups with less than 100 employees need to optimize the use of LLMs within constraints of limited computational resources and workforce. These industries should emphasize prompt engineering and transfer learning techniques to utilize robust small LLMs with up to 3 billion parameters with permissive licenses. Further, these industries should focus on monolingual uni-modal tasks and actively perform the inference on a few hundred samples. To reduce the inference duration, these industries should opt for optimization techniques such as quantization. Moreover, these industries encounter challenges such as potential reductions in model accuracy, cost considerations, and the need for technical expertise, some of these can be addressed by partnering with AI consulting firms.
2) Medium-scale industries up to 1000 employees should focus on utilizing the RAG-based pipelines and domain-specific parameter efficient fine-tuning and distillation techniques for LLMs up to 7B parameters. Additionally, these industries can develop domain-specific adapters to enhance LLMs’ performance on specific tasks. These industries can explore moderate multi-modal (text + vision) tasks. Additionally, the key challenges for medium-scale industries are scalability and domain adoption.
3) Large scale industries such as MNCs should focus on continuous pre-training of LLMs while ensuring compliance with regulatory requirements. These industries can leverage LLMs effectively across multi-lingual, cross-lingual, and multi-modal generation tasks. Training of such models can take from a few weeks to months, which requires high-quality data and huge compute as well. These industries should focus on establishing several collaborative tools and frameworks to maximize LLM utilization.
For all industries, we recommend using open-source models with proper licenses to address ethical concerns and comply with LLM regulatory guidelines. Additionally, robust testing and validation protocols are essential to meet industry standards. Fostering strong collaborations and knowledge sharing between industry and academia is crucial for advancing responsible LLM development and deployment.
6 Conclusions
This study delves into the utilization of Large Language Models (LLMs) through an industrial lens, with a specific focus on identifying roadblocks to their adoption. It meticulously examines various pitfalls and provides potential solutions. Moreover, this study offers a guide to organizations of all sizes to maximize the utilization of LLMs for industrial use cases. By identifying pitfalls and suggesting potential directions, the study offers a strategic roadmap for optimizing LLM effectiveness in industrial operations. We plan to maintain a GitHub repository with the most recent papers in the field.
7 Limitations
Our study has the following limitations.
Scope. To provide a practical guide to various industries, we restrict our scope to only Natural Language Generation (NLG) applications. Prospective works should focus on providing an extensive guide to various other tasks as well.
Coverage. With the rapid development of LLMs and the voluminous research in this field, it’s not feasible to comprehensively cover all the papers. Recognizing this, our survey has focused specifically on industry-related papers. This allowed us to delve deeper and gain an understanding of the unique requirements and challenges faced within industrial applications of LLMs.
Confidentiality. Due to the confidential nature of the industrial applications not many details were available for specific scenarios or challenges. Hence, we only focused on providing recommendations/insights that can be applicable to a broad range of industrial applications.
8 Ethics Statement
To our knowledge, this study presents minimal ethical concerns. However, to maintain transparency, we provide a detailed analysis of all 68 papers present in the survey in Section C. Each paper is reviewed by at least three individuals to validate its claims and findings. We conduct the industrial case study, following the guidelines outlined by the ACL ethics review policy  [^2] https://aclrollingreview.org/ethicsreviewertutorial, thereby Ethics Review Boards (ERB) approval is not necessary. It’s important to note that our research involving human subjects does not entail the collection of any medical or sensitive information from the users.
Appendix A Survey Papers Selection Criteria
We used keywords such as “large language models”, “LLM” and “LLMs for industrial applications” for selecting the relevant papers. We selected the majority of papers from the reputed databases including the ACL Anthology [^3] https://aclanthology.org/, ACM Digital library [^4] https://dl.acm.org/, Google Scholar [^5] https://scholar.google.com/, which are known for hosting peer-reviewed articles that meet high academic standards. Subsequently, we finalize suitable research papers for the survey based on the following criteria.
Criteria Number of papers
arXiv version 37
Non organizational papers 10
Not related to application 6
Relevant 68
Total 121
Table 2: Survey papers filtration criteria.
  • The paper should be a peer-reviewed and published version.
  • At least one of the paper’s authors should be from the industry.
  • Paper should use at least one or more LLM.
  • The paper should report at least one real-world application using LLM(s).
Necessary Concessions: We believe that having at least one author from the industry brought the following advantages.
  • We found that considering papers with only researchers from industry led to very few research papers. Also, in recent times, collaboration between academia and industry has rightfully expanded resulting in more practical and applicable research works.
  • Also, they brought practical perspectives that were grounded in real-world applications and challenges.
In total, we have collected 121 research papers, and out of them, we have discarded 53 that do not fall under one or more above-mentioned criteria as mentioned in Table 2. We have omitted 40 papers because they are not peer-reviewed and 10 more papers came from the non-organizations typically submitted by academic labs/universities. Moreover, we have discarded six papers, which did not discuss any industrial application. After applying the filtering criteria we left with 68 relevant papers. This distribution of the list of papers from various industrial organizations is mentioned in Figure 3.
MicrosoftGoogleAmazonIBMNAVER AIOthers5555101010101515151520202020212121211010101077775555333313131313Frequency
Figure 3: Distribution of research papers from industrial organizations. Others include Apple, Sony, Alibaba, Allen Inst for AI, JP Morgan, Nvidia, Adobe.
Appendix B Industrial Case Study Details
1. Participant level of expertise in LLMs? $\square$ Beginner $\square$ Intermediate $\square$ Proficient $\square$ Expert $\square$ NA 2. Application Domain $\square$ Healthcare $\square$ Banking $\square$ Financial $\square$ Retail $\square$ Security $\square$ Privacy $\square$ Legal $\square$ Marketing & Advertising $\square$ Education $\square$ Media and entertainment $\square$ Human Resources(HR) $\square$ eCommerce $\square$ Other:   3. What is the name of task that LLM(s) performs in your project? 4. Type of data used? $\square$ Tabular $\square$ Image $\square$ Video $\square$ Audio $\square$ Text $\square$ More than one modality $\square$ Other:   5. How are the LLMs used? $\square$ Fine-tuning $\square$ Zero-shot or Few-shot $\square$ In-context learning $\square$ Other:
Table 3: Questionnaire for industrial case study: Part 1
6. Did you consider any of the following Trust attributes or guard rails while designing/implementing the LLM-based solution? $\square$ Security $\square$ Robustness $\square$ Privacy $\square$ Bias & Fairness $\square$ Interpretability or Explainability $\square$ Toxicity $\square$ Hallucination $\square$ None $\square$ Other:   7. Name of the LLMs being used? $\square$ LLaMA $\square$ LLaMA-2 $\square$ Falcon $\square$ Mistral $\square$ GPT3.5 (ChatGPT) $\square$ GPT4 $\square$ MPT $\square$ Meta OPT $\square$ Bard $\square$ PaLM $\square$ Pythia $\square$ Cerebras-GPT $\square$ NA $\square$ Other:   8. What are the risks associated with the LLMs being used in your project? $\square$ Security and Safety $\square$ Reputation $\square$ Quality of service $\square$ Revenue $\square$ License $\square$ NA $\square$ Other:
Table 4: Questionnaire for industrial case study: Part 2
We have created a questionnaire to conduct the industrial case study as shown in Fig 3.
Appendix C Survey Papers Checklist
This paper provides a review of 68 papers and for each paper, we reported 22 features as mentioned in Table 5. We briefly describe each feature in the master table for better understanding.
  • Paper: Citation of the paper.
  • Venue: The venue where the paper was published.
  • Year: Year of paper publication.
  • LLM name: Names of the LLMs used in the paper.
  • Organization: Name of the industrial organization involved in the work.
  • Domain: Domain information of the application in the paper.
  • Application: The type of application under which the work was categorized into.
  • Use case: The information of how the paper leverages an LLM in a specific scenario or a task.
  • Dataset Name: Datasets used by the paper for modeling and evaluation.
  • Prompting Strategy: Prompting strategies used in the paper.
  • Evaluation metrics: Details of the evaluation metrics used in the paper.
  • Application life cycle: Information of application’s life cycle stage.
  • GitHub: Link to the GitHub repository, if any, that was published in the paper.
  • License: This field indicates if the paper contains license-related information.
  • Privacy:This field indicates if the paper contains privacy-related information.
  • Use cases: This field indicates if the paper mentions a use case or not.
  • Limitations: Major limitations of the paper, if any.
[…]
Paper Venue Year LLMs used Organization Domain Application Use case Dataset Name Prompting strategy Evaluation metrics Application life cycle Github License Privacy Use cases Limitations
1 Li et al. (2023c) EMNLP Industry Track 2023
ChatGPT, GPT-4, BloombergGPT,
GPT-NeoX, OPT66B, BLOOM176B,
FinBERT
J.P. Morgan AI Research Financial Analytics Financial text analysis
FPB/FiQA/TweetFinSent, Headline,
NER, REFinD, FinQA/ConvFinQA
Zero-shot, Few-shot and CoT Accuracy, F1 Score Conceptualization/PoC NA Yes NA Yes NA
2 Chen et al. (2023) EuroSys 2024 GPT-3.5, GPT-4 Microsoft Fault diagnosis Cloud management Cloud incident root cause analysis
653 incidents from Microsoft’s transport service
Zero-shot Micro and Macro F1-score Conceptualization/PoC NA NA NA NA Effectivness of the method depends on incident monitors/alerts.
3 Bairi et al. (2023) FMDM@NeurIPS 2023 GPT-4-32k Microsoft research Software Code generation Automate repository level code planning tasks Proprietary Zero-shot Block metrics, Edit metrics Conceptualization/PoC NA NA NA NA
Dynamic languages may not be ideal for a coded plan approach
4 Mani et al. (2023) HotNet’s 2023
GPT-4, GPT-3, Text-davinci-003,
Bard
[…]
High quality domain specific code synthesis is still an open challenge
5 Ding et al. (2023) ACL 2023
CodeGen-350M,CodeGen-2B,
CodeGen-6B,CodeGen-16B,
AWS AI Labs Software Code generation Static code analysis for completion Function completion dataset None
Percentages of AST errors,
Undefine variable, unused variables etc.
Conceptualization/PoC NA NA No Yes
Cross-file context based broader categorization of errors
was not employed
6 Li et al. (2023a) UIST 2023
GPT-3.5
Legacy (text-davinci-003),
Legacy (code-davinci-002),
Legacy (text-davinci-edit-001)
NA Software Code generation Webpage customization using LLMs NA Fewshot (In-context learning) NA Prototype NA No No Yes
Poor performance on complex website customization
7 Phung et al. (2023) ICER 2023 GPT-3.5, GPT-4 Microsoft Software Code generation
LLM tutor for programming education
NA Zero-shot Match Conceptualization/PoC NA NA NA Yes
Limited to python language and introductory educational content
8 Gupta et al. (2023) ESEC/FSE 2023 CODEX, CODEt5 Microsoft Software Code generation
Prediction of code edits using LLMs
C3PO, Overwatch Fewshot (In-context learning) Exact match of the code segment Conceptualization/PoC NA NA NA Yes
Proposed approach may fail due to LLMs hallucination
and context length requirements
9 Dinh et al. (2023) NeurIPS 2023 CODEGEN, INCODER AWS Software Code generation
Context based code completion
[…]
Proposed method may not be aligned to general software development
setting as buggy datasets are based on programming contest submissions
[…]
Skewed age distribution in pilot subjects and pilot deployment time
12 Valencia et al. (2023) CHI 2023 NA Google Research Accessibility Conversational Evaluation of LLMs as a tool for AAC users Proprietary None NA Conceptualization/PoC NA NA NA Yes NA
13 Gadiraju et al. (2023) FAccT 2023 LaMDA Google Research Accessibility Conversational Chatbot NA None Human evaluation NA NA Yes No Yes Limited diversity of identity within individual focus groups
14 Feng et al. (2023) NeurIPS 2023
Codex, GPT-3.5, GPT-3.5-chat,
GPT-4
Google Generic Data Generation
Visual planning for text-to-image generation
NSR-1K, 3D-FRONT Fewshot (In-context learning)
CLIP cosine similarity,
GLIP accuracy,
Attribute binding Accuracy,
KL divergence
[…]
Generation of overly dense layouts and unusual sized bounding boxes
15 Golde et al. (2023) ACL 2023
Used existing LLMs from Hugginhface,
openAI, Azure, Anthropic, Cohere
[…]
Evaluation cover only subset of commonly encountered tasks
16 Yu et al. (2023b) EMNLP Industry Track 2023 GPT-4, LLaMA Amazon Financial Forecasting
Explainable financial time series forecasting
Stock price data, Company profile data,
Finance/Economy News Data
Zero-shot and Few-shot
Binary Precision, Bin Precision,
MSE, ROUGE-1,2
Conceptualization/PoC NA Yes Yes NA
Generalization to other types of financial temporal data unexplored
17 De La Torre et al. (2023) NeurIPS 2023
Language model for mixed reality (LLMR)
Dall.E-2
GPT-4
Microsoft, Microsoft Research Generic Frameworks
Generation of interactive 3D objects
NA NA
Error rate, Average generation time.
Development NA NA NA Yes
For complex tasks manual code editing might be still necessary
18 Wang et al. (2023a) CHI 2023 PaLM Google Research NLP Frameworks
Conversational Interaction
with Mobile UI
PixelHelp , AndroidHowTo,
Rico, Screen2Words,
Zero-shot and Few-shot
Grammar Correctness, UI Relevance,
Question Coverage, BLEU, CIDEr,
ROUGE-L, and METEOR, Micro-F1
[…]
Fails to handle generation of incorrect or irrelevant information
19 Hoshi et al. (2023)
EMNLP System
demonstrations
2023
Llama-2 Chat (13B, 70B),
WizardVicunaLM-13B,
Vicuna
Kioxia Corporation Generic Frameworks
Framework for knowledge intensive tasks
KILT Benchmark NA
Exact Match, F1, Accuracy, RL,
R-precision
[…]
20 Zhang et al. (2023) CoRL 2023 LLaMA-13b Google AI Robotics Frameworks
LLM guided skill chaining
ALFRED Zero-shot NA Development NA NA No Yes
Greedy skill chaining may not be optimal for consistent behaviour generation
21 Liu et al. (2023b) EMNLP 2023
GPT-3.5-turbo, text-davinci-003,
GPT-4-32k
Microsoft health futures Healthcare Frameworks
Evaluation of GPT-4 on understanding and
generation of radiology tasks
MS-CXR-T, RadNLI, Chest ImaGenome,
MIMIC, Open-i
Zero-shot, Few-shot, One-shot,
Many-shot, CoT
macro F1, micro F1, RougeL,
CheXbert score
Deployment NA NA Yes Yes
Qualitative evaluation of the findings on summarization task is
limited to a single radiologist
22 Jiang et al. (2023b) EMNLP 2023 GPT-3.5-Turbo-0301 and Claude-v1.3 Microsoft NA None
Prompt compression for higher inference speed
GSM8K, BBH, ShareGPT,
Arxiv-March23
[…]
Performance drops when compression of prompts go beyond 25%
23 Yang et al. (2023) EMNLP Industry Track 2023 GPT-4, GPT3.5, LLaMA-2 Microsoft NLP Question-answering
Domain specific industrial QA
MSQA Zero-shot
BLEU, ROUGE, METEOR, BERTScore,
F1, Key-word/Span-Hit-Rate (KHR),
Can-Answer-Rate (CAR), LLM-based Metrics
Conceptualization/PoC NA NA NA NA
Only works with English data
24 Zeng et al. (2023) ICAIF 2023 GPT-3.5-turbo J. P. Morgan AI Research Finance Question-answering
Dynamic workflow generation
NCEN-QA, NCEN-QA-Easy,
NCEN-QA-Intermediate, NCEN-QA-Hard
Zero-shot NA Prototype NA NA Yes Yes NA
25 Hu et al. (2023) ICCV 2023 GPT-3 Microsoft, Allen Institute of AI Vision Question-answering
Natural language based question-aware caption model
[…]
Focuses only on knowledge-based VQA tasks
26 Liu et al. (2023a) NeurIPS 2023
GPT-3.5 turbo, GPT-4,
ChatGLM, LLaMA,
Vicuna, Alpaca
Alibaba Group, Ant Group Medical Question-answering
Medical domain QA
[…]
Excluding non-textual questions might introduce unexpected bias
27 Imani et al. (2023) EMNLP Industry Track 2023 text-davinci-002, PaLM Microsoft NLP Reasoning Mathematical Reasoning MultiArith dataset Zero-shot, Few-shot and CoT Accuracy Conceptualization/PoC NA Yes NA NA
Non trivial probability of producing incorrect results
using algebraic and pythonic expressions
[…]
29 Li et al. (2023b) CIKM 2023
GPT-Neo-1.3B,
GPT-Neo-2.7B,
GPT-J-6B,
Falcon-7B-Instruct
Amazon Alexa AI Political, Education Reasoning
Steerability of LLM based on persona
OpinionQA None
User study
Conceptualization/PoC NA NA No Yes Complex personas may not be possible
30 Ye et al. (2023) SIGIR 2023 CODEX Alibaba Group Generic Reasoning
Reasoning on large tables based on textual prompts
TabFact, WikiTableQuestion, FetaQA Fewshot (In-context learning)
binary classification accuracy,
denotation accuracy, BLEU,
ROUGE-1, ROUGE-2 and ROUGE-L
Conceptualization/PoC NA Yes Yes Yes NA
31 Pawlowski et al. (2023) ICML Workshop 2023 GPT-3.5, GPT-4 Microsoft Research Generic Reasoning
LLM based causal QA system
NA Text completion NA Conceptualization/PoC NA NA NA Yes Performance decreases with increased context length
32 Sanner et al. (2023) RecSys 2023 PaLM Google Retail Recommended systems
Visually augmented real-time conversations
Proprietary Completion, zero-shot and Few-shot Mean NDCG Conceptualization/PoC NA No Yes Yes NA
33 Yamazaki et al. (2023) Advanced robotics 2023 Hyperclova LINE corporation Generic Recommender systems Voice Chatbot Proprietary Fewshot (In-context learning)
Technical score on Informativeness,
Naturalness, Likability,
Satisfaction with dialog
Testing NA NA No Yes
For low frequency words gives long
responses which floods user with information and hallucination
34 Maragheh et al. (2023) ICML Workshop 2023 PaLM2 Walmart Global Tech Retail Recommender systems
Enhance the capabilities of recommendation
systems
Proprietary Zero-shot MRR, NDCG Conceptualization/PoC NA NA NA NA NA
35 Alfassy et al. (2022) NeurIPS 2022 CLIP
IBM Research,
MIT-IBM AI-Watson Lab
Vision Retrieval
Evaluation of LLMs on expert tasks for \
image-to-text and text-to-image retrieval
FETA
Zero-shot, Few-shot, One-shot,
Many-shot
Accuracy Conceptualization/PoC NA NA NA NA
FETA contains only a small subset of available
technical documents for different expert V&L data domains
[…]
Limited ability to update knowledge to new domains
37 Wang et al. (2023b) EMNLP 2023
Text-davinci-001, Text-davinci-003,
GPT-4, Babbage, curie
Microsoft Research Generic Retrieval Query expansion based retrieval systems MS-MARCO, TREC DL 2019 Fewshot MRR, nDCG Conceptualization/PoC NA NA NA Efficiency of retrieval system
38 Liu et al. (2023c) CHI 2023 GPT3 Google Research Generic Retrieval
Augmenting video conferencing
with visual captions
[…]
Visual captions in conversations should have a threshold
to filter out potentially distracting or inappropriate content
39 Lu et al. (2023b) NeurIPS 2023
GPT2-S (117M),
GPT2-L (774M) [29],
OpenLLaMA-7B (7B)
AWS GAIIC,
AWS AI
Healthcare Retrieval Writing radiology reports from medical images MIMIC-CXR None
Factual completeness and correctness
F1-CXB-14 score, F1-CXB-5,
BLEU4, ROUGE-L
[…]
soft visual prompt doesn’t receive consistent attention,
especially when using LLMs.
40 Alaofi et al. (2023) SIGIR 2023 text-davinci-003 Microsoft Generic Retrieval
Generation of query variants for building
test collections and document pool
UQV100 One-shot
Jaccard Index, RBP, RBO
Conceptualization/PoC NA NA NA Yes
Small size human-generated data is not
sufficient for few-shot prompting
41 Park and You (2023) EMNLP Industry Track 2023 CTI-BERT IBM T. J. Watson Research Center Security Security Cyber threat intelligence
Attack description, Security Textbook,
Academic Paper, Security Wiki,
Threat reports, Vulnerability
NA Micro and Macro F1 Score Conceptualization/PoC NA Yes Yes Yes Pretrained only on English data.
42 Foley et al. (2023) ACL 2023
BERT, GPT, BLOOM, codegen-350M,
DialoGPT, DistilGPT2,
OPT, GPT-Neo, xlnet-base-cased,
multilingual-miniLM-L12-v2
IBM Research Generic Security
Tracing back to the origin of fine-tuned models to
alleviates the problem of accountability of LLMs
GitHub, The BigScience ROOTS Corpus,
CC-100, Reddit, and THEPILE
NA F1, ROC Conceptualization/PoC NA Yes Yes Yes
Considered only a limited number of LLMs for the study.
43 Peng et al. (2023) ACL 2023 text-embedding-ada-002, BERT
Microsoft Research Asia, Sony AI,
Micorsoft STC Asia
Security Security
Copy right protection of EaaS (Embeddings as
a Service) LLMs
SST2, Mind, Enron Spam, AG news NA Accuracy, Detection performance Conceptualization/PoC NA NA NA NA NA
44 Deng et al. (2023) WWW 2023 GPT-3, PaLM Google Research Finance Sentiment Analysis
Labels generation for financial data
FiQA-News Fewshot (In-context learning) Accuracy Conceptualization/PoC NA NA No Yes -
[…]
The performance of the filter models are not very competitive
46 Sun et al. (2023) EMNLP Industry Track 2023
GPT-3.5-turbo-0301, Falcon 40B-instruct,
Falcon 7B-instruct, Dolly-v2-12b
Apple Security Societal Impact
Comprehensive handling of controversial issues
DELPHI Zero-shot
Controversy Acknowledgement Rate,
Comprehensiveness Answer Rate
[…]
Dataset may not cover all the controversial questions.
and may contain expired ground truth controversy labels
47 Athiwaratkun et al. (2023) ICLR 2023 Decoder-only transformer model AWS AI Labs Software Code generation
Evaluation of LLMs on multilingual
programming datasets
[…]
Does not support language-specific functionalities
48 Laskar et al. (2023) EMNLP Industry Track 2023
GPT-4, GPT3.5, PaLM-2,
LLaMA-2 13b, 7b
Dialpad Canada Inc NLP Summarization Business meeting summarization QMSUM, AMI, ICSI Zero-shot ROUGE, BERTScore Conceptualization/PoC NA Yes Yes Yes
Generalizability to domain-specific datasets is in question
because only academic datasets were used for testing
49 Fetahu et al. (2023) EMNLP Industry Track 2023 FLAT-T5 Amazon NLP Summarization Summarization of length product titles NA NA ROUGE, BLEU Conceptualization/PoC NA NA NA NA Do not guarantee inclusion of salient words in the summary
50 Acharya et al. (2023) RecSys 2023 Alpaca-LoRa Sony Research India Retail Summarization
Generation of product descriptions sans webscrapping
MovieLens, Goodreads Book graph Fewshot (In-context learning)
Hit Rate, Normalized Discount
Cumulative Gain (NDCG),
Mean Reciprocal Rank (MRR)
Conceptualization/PoC NA NA NA NA
Generates factually incorrect descriptions
51 Jin et al. (2023) ESEC/FSE 2023 GPT-3.X Microsoft Infrastructure Summarization
Cloud outage management
Proprietary NA BLEU-4, ROUGE-L and METEOR Deployment NA NA NA Yes Evaluation metrics not fully reflect readability and usefulness of outage summary
52 Shen et al. (2023) CHI In2Writing Workshop 2023 NA Allen Institute of AI NLP Summarization
Evidence based knowledge generation
NA Fewshot (In-context learning) NA NA NA NA NA Yes NA
53 Zhao et al. (2023) EMNLP Industry Track 2023
GPT4, TULU, Pythia, Alpaca,
Vicuna, LLaMA-2, GPT-3.5
Allen Institute for AI NLP Table-to-text-generation
Evaluation of LLMs on table-to-text generation
LOTNLG, F2WTQ Zero-shot and Few-shot
BLEU, ROUGE, SP-Acc, NLI-Acc,
TAPAS-Acc, TAPEX-Acc, Exact-match,
F1 Score, Accuracy
Conceptualization/PoC NA Yes NA Yes NA
54 Mihindukulasooriya et al. (2023) ISWC 2023 GPT4, Llama2, FLAN-T5 IBM Research NLP Table-to-text-generation
Automation of meta data
generation and enrichment
NA None NA Conceptualization/PoC NA NA NA NA NA
55 Singha et al. (2023) NeurIPS 2023 GPT-3.5 (text-davinci-003 endpoint) Microsoft Generic Table-to-text-generation
Noise induction to better
understanding of table structures
AirQuality, HousingData, Diabetes,
Wine Testing, Iris, Titanic,
and ENB2012_data
[…]
Performance of structural tasks with downstream task such as table
question answering remains an open challenge.
56 Wake et al. (2023) IEEE Access 2023 ChatGPT Microsoft Robotics Task Planning
Translating natural language instructions
to executable robot actions
[…]
57 Saadany and Orasan (2023) EMNLP Industry Track 2023 GPT- text-embedding-ada-002 Kingfisher Labs Ltd , Just Access Legal Tool
Automatic linkage of judgements to bookmarks
in court hearing videos
UK National Archive Zero-shot Mean Average Precision (MAP), Recall Conceptualization/PoC NA Yes Yes NA NA
58 Petridis et al. (2023) CHI Extended Abstract 2023 - Google Research Generic (HCI) Tool input-output interaction , Frame change NA Zero-shot Questionnaire Conceptualization/PoC NA NA NA Yes
Needs formal evaluation and in-depth analysis on
how functional prompts effect prototyping process
59 Dibia (2023) ACL 2023 NA Microsoft Research Generic Tool
Automatic generation of grammar-agnostic
visualizations and infographics
Proprietary Zero-shot and Few-shot
Visualization Error Rate (VER),
Self-Evaluated Visualization Quality
(SEVQ), code accuracy,
data transformation, goal compliance,
visualization type, data encoding,
and aesthetics
[…]
Code execution step increases computational complexity.
60 Singh et al. (2023) ICRA 2023
text-davinci-*,
Codex,
GPT3
Nvidia corporation Robotics Tool Generate programmatic robot instructions using LLMs NA Fewshot (In-context learning)
Success rate(SR),
Goal condition recall(GCR),
Executability(Exec)
[…]
Robotic action success feedback is not
shared with agent leading to failure scenarios
61 Awasthi et al. (2023) EACL 2023
mT5-Large,
PaLM
Google research India NLP Translation Translating English datasets to several other languages MTOP, MASSIVE Zero-shot and Few-shot corrected Exact Match Prototype NA NA No Yes
Computationally expensive
[…]
The tool’s effectiveness not tested for decoder-only models.
63 Ramakrishna et al. (2023) EMNLP 2023
GPTNeo-2.7B, GPTJ-6B,
Open-LLaMA-7B,
RedPajama-7B,GPT3.5-Turbo,
GPT4
Amazon Alexa AI Security Trustworthy AI Evaluating the LLMs for Hallucinations DBpedia, TriviaQA Zero-shot
BLEU, ROUGE, METEOR,
BERTScore, AlignScore
[…]
Lack of diversity in test set.
64 Kotek et al. (2023) CI 2023 Not disclosed Apple Generic Trustworthy AI Identify the presence of gender bias in LLMs Proprietary Zero-shot NA Conceptualization/PoC NA NA NA Yes Results may not reflect the real gender bias
[…]
Only consider binary gender when analyzing biases
66 Kim et al. (2023) NeurIPS 2023
OPT-350M
OPT-1.3B
OPT- 2.7B
NAVER AI Lab, Parameter Lab Generic Trustworthy AI Probing for PII in a given LLM Pile Fewshot (In-context learning)
Sentence matching,
Likelihood ratio
Conceptualization/PoC NA NA Yes/ Yes
Evaluation dataset uses private
information sourced from open-source datasets
67 Rebedea et al. (2023)
EMNLP System
demonstrations
2023
text-davinci-003, GPT-3.5-turbo,
falcon-7b-instruct, llama2-13b-chat
NVIDIA Generic Trustworthy AI
Tool kit for adding programmable guardrails for
conversational LLMs
Anthropic Red-Teaming and
Helpful datasets
[…]
Toolkit not suitable as standalone solution
68 Candel et al. (2023)
EMNLP System
demonstrations
[…]
Deploy and test efficiency of wide variety of LLMs
on private databases and documents
[…]
Datasets, Biases and Offensiveness, Usage, Carbon footprint,
Hallucinations of LLMs
Table 5: Master table of the survey with 68 research papers.